OpenStudy (anonymous):

Write the equation of the line in slope-intercept form with a slope of -5 and a y-intercept of -1.

OpenStudy (anonymous):

y=-5x-1

OpenStudy (anonymous):

y=mx+c where m is the slope and c the y intercept

OpenStudy (ghazi):

first of all write y=mx+c now put the value of slope in the equation y=-5x+c now you've been given y intercept which means put x=0 so y=-1, when x=0 -1=5*0+c therefore c=-1 and your equation is y=-5x-1 hope that helps @ambermarie

OpenStudy (anonymous):

Thank you :D